1+ 1 over x = 30 over x squared
Answer by checkley71(8403)

1+1/X=30/X~2 OR (X+1)/X=30/X~2 OR 30X=X~3+X~2 OR DIVIDING ALL TERMS BY X WE GET
30=X~2+X OR X~2+X-30 OR (X-5)(X+6)=0 OR X=5 & X=-6
PROOF
1+1/5=30/25 OR 6/5=30/25 OR 6/5=6/5
1+1/-6=30/(-6)~2 OR 5/6=30/36 OR 5/6=5/6